OpenStudy (anonymous):

how many number of ways in which each committee can be selected..8 people from a group of 12

Parth (parthkohli):

More like choosing 8 from 12. Which makes \(12C8\).

OpenStudy (anonymous):

12!/8!

Parth (parthkohli):

\[12C8 = {12! \over (12 - 8)!8!} \]

Parth (parthkohli):

Manage the rest, okay?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

but..wouldnt it be 12!/(12-8)!4!??

OpenStudy (amistre64):

no

OpenStudy (amistre64):

\[nPr=\frac{n!}{(n-r)!}\] \[nCr = \frac{nPr}{r!}=\frac{\frac{n!}{(n-r)!}}{r!}\to\ \frac{n!}{(n-r)!\ r!}\]